what is the equation of a line through the point ( -2, 3 ) and a slope of 4
Found 2 solutions by ankor@dixie-net.com, ikleyn:
Answer by ankor@dixie-net.com(22740)

what is the equation of a line through the point ( -2, 3 ) and a slope of 4
:
x=-2, y=3, m=4
Using the slope/intercept form mx + b = y, replace x, y, and m
4(-2) + b = 3
-8 + b = 3
b = 3 + 8
b = 11
the equation
y = 4x + 11

Answer by ikleyn(52781)

.

If you want to learn the subject from  "a"  to  "Z"  and to learn on how to solve such problems  on your own,  look into the lesson
    - Equation for a straight line having a given slope and passing through a given point
in this site.

My other lessons in this site related to the slope of a straight line are

    - Find the slope of a straight line in a coordinate plane passing through two given points
    - Equation for a straight line having a given slope and passing through a given point
    - Solving problems related to the slope of a straight line
    - Equation for a straight line in a coordinate plane passing through two given points
    - Equation for a straight line parallel to a given line and passing through a given point
    - Equation for a straight line perpendicular to a given line and passing through a given point

    - OVERVIEW of lessons related to the slope of a straight line

Consider these lessons as your textbook,  handbook,  tutorials and  (free of charge)  home teacher.